# scSiameseClu: A Siamese Clustering Framework for Interpreting single-cell RNA Sequencing Data

### v1

@@ Abstract
22footnotetext: The code is available at the link: https://github.com/XPgogogo/scSiameseClu.
@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) technology represents a significant advancement in bioinformatics, enabling the capture of comprehensive genetic information from individual cells Shapiro et al. (2013); Yang et al. (2024). Cell clustering, a key step in single-cell RNA sequencing analysis, groups cells by their gene expression patterns to uncover the complex characteristics of distinct cell populations and provide insights into their biological functions and interactions Kiselev et al. (2019). Moreover, clustering gene expression patterns further contributes to various downstream tasks, such as marker gene identification and cell type annotation Haghverdi et al. (2016). Cluster analysis of scRNA-seq data has been a vibrant research area over the past decade.
In recent years, advanced computational methods have been increasingly explored to address the challenges of analyzing high-dimensional and sparse scRNA-seq data. Classical clustering methods, such as K-means, are straightforward and computationally efficient but struggle to capture the complex, nonlinear relationships in scRNA-seq data. Meanwhile, recent research has focused on applying deep learning frameworks Ning et al. (2024) for learning representations of scRNA-seq data and classifying samples into distinct clusters. A commonly used method is self-supervised learning methods, which uncover effective representations of scRNA-seq data by reconstructing the original input data Eraslan et al. (2019); Lopez et al. (2018); Tian et al. (2019). However, these methods primarily focus on extracting features from individual cells, overlooking the complex relationships between cells that are essential for understanding cellular diversity. Benefiting from the powerful utilization of graph information Ning et al. (2021), graph neural networks (GNNs) have been applied to analyze scRNA-seq data by modeling cells as nodes and their interactions as edges, effectively capturing both gene expressions and cell graphs to comprehensively represent cellular heterogeneity Luo et al. (2021); Wang et al. (2021b); Gan et al. (2022); Zhan et al. (2023).
@@ 2 Related work
2.1 Deep Clustering for scRNA-seq.
In scRNA-seq data analysis, early clustering algorithms like Phenograph Levine et al. (2015), MAGIC Van Dijk et al. (2018), and Seurat Wang et al. (2025) utilize k-nearest neighbor (KNN) graphs to model cell relationships, while SIMLR Wang et al. (2018) and MPSSC Park and Zhao (2018) employ multiple kernel functions to derive robust similarity measures from different data representations. These methods, however, often struggle with high-dimensional data and complex nonlinear features, making them susceptible to noise. Deep learning approaches have since become prominent. DCA Eraslan et al. (2019) employs a Zero-Inflated Negative Binomial (ZINB) autoencoder to model scRNA-seq data distributions, capturing nonlinear gene dependencies. SCVI Lopez et al. (2018) and SCVIS Ding et al. (2018), which rely on autoencoders for dimensionality reduction, often face over-regularization due to their Gaussian distribution assumption. Additionally, scDeepCluster Tian et al. (2019) combines ZINB-based autoencoders with deep embedding clustering to optimize feature learning and clustering. Most approaches ignore inter-gene and inter-cell correlations, focusing solely on individual cell expression profiles.
2.2 Graph Clustering for scRNA-seq.
Recently, graph-based clustering methods have gained considerable attention for their effectiveness Tu et al. (2021); Liu et al. (2022); Ning et al. (2025). Methods like scGAE Luo et al. (2021) and graph-sc Ciortan and Defrance (2022) utilize graph autoencoders to embed scRNA-seq data while preserving topological structures. scGAC Cheng and Ma (2022) constructs cell graphs and employs a self-optimization approach for simultaneous representation learning and clustering. Notably, scGNN Wang et al. (2021b) uses GNNs and multi-modal autoencoders to aggregate cell-cell relationships and model gene expression patterns, while scDSC Gan et al. (2022) combines a ZINB model-based autoencoder with GNN modules using a mutually supervised strategy. These methods often struggle with representation collapse, limiting their ability to model cell relationships.
@@ 3.1 Problem Formulation
3.3 Dual Augmentation Module.
@@
3.4 Siamese Fusion Module.
@@
Latent Correlation Refinement. Similarly, latent correlation refinement (LCR) optimizes relationships between latent embeddings across different augmented views by aligning corresponding embeddings. First, we project gene expression embeddings $\mathbf{Z}^{G1}$ and $\mathbf{Z}^{G2}$ into $\widetilde{\mathbf{Z}}^{G1}$ and $\widetilde{\mathbf{Z}}^{G2}\in\mathbb{R}^{d\times K}$ with using a readout function $\mathcal{R}(\cdot):\mathbb{R}^{d\times N}\rightarrow\mathbb{R}^{d\times K}$ , formulated as $\widetilde{\mathbf{Z}}^{G1}=\mathcal{R}\left(\left(\mathbf{Z}^{G1}\right)^{\mathrm{T}}\right)$, $\widetilde{\mathbf{Z}}^{G1}=\mathcal{R}\left(\left(\mathbf{Z}^{G2}\right)^{\mathrm{T}}\right)$. Likewise, we can project cell graph embeddings $\mathbf{Z}^{C1}$ and $\mathbf{Z}^{C2}$ into $\widetilde{\mathbf{Z}}^{C1}=\mathcal{R}\left(\left(\mathbf{Z}^{C1}\right)^{\mathrm{T}}\right)$ and $\widetilde{\mathbf{Z}}^{C2}=\mathcal{R}\left(\left(\mathbf{Z}^{C2}\right)^{\mathrm{T}}\right)$.
@@
3.5 Optimal Transport Clustering.
@@
In this context, we regard the target distribution $P$ as the transport plan matrix derived from optimal transport theory, while $-\text{log}Q$ serves as the cost matrix. We impose the constraint $P^{T}\bm{1}_{N}=N\bm{\pi}$, where $\pi$ indicates the proportion of cells assigned to each cluster, estimated from intermediate clustering results. Thus, we can ensure that the resulting cluster distribution aligns with the defined mixing proportions. Given the substantial computational expense associated with direct optimization techniques, we utilize the Sinkhorn distance Sinkhorn (1967) to facilitate quicker optimization through an entropic constraint. The optimization problem, which includes a Lagrange multiplier for the entropy constraint is formulated as:
@@ 4 Experiments
In this section, we first validate our method through extensive experiments, demonstrating its superior performance in scRNA-seq clustering. We then show that the embeddings generated by our method effectively mitigate the embedding collapse issue in graph-based models. We evaluate its utility in downstream tasks, like cell type annotation, and perform ablation studies to assess the contributions of individual components. Additional details are in the supplementary material.
4.1 Experimental Setup.
@@
Evaluation Metrics. To illustrate the effectiveness of our approach, we evaluate clustering performance through three widely recognized metrics: Accuracy (ACC), Normalized Mutual Information (NMI) Strehl and Ghosh (2002), and Adjusted Rand Index (ARI) Vinh et al. (2009). Higher values for these metrics indicate better clustering results.
@@
4.4 Ablation Study
(a) Major components evaluaion
(b) Component Effects in SFM
Figure 7: Ablation study on  Shekhar mouse retina cells dataset.
@@ 5 Conclusion
6 Acknowledgements
This work is partially supported by the National Natural Science Foundation of China (Grant Nos. 92470204 and 62406306), the National Key Research and Development Program of China (Grant No. 2024YFF0729201), and the Science and Technology Development Fund (FDCT), Macau SAR (file no. 0123/2023/RIA2, 001/2024/SKL).
@@ C.3 Gene expression and Cell Graph Autoencoders
[…]
@@
[…]
@@ D.1 Evaluation Metrics
Macro $\text{Macro Precision}=\frac{1}{N}\sum^{N}_{i=1}\frac{TP_{i}}{TP_{i}+FP_{i}}$ (27)
@@
Macro $\text{Macro Recall}=\frac{1}{N}\sum^{N}_{i=1}\frac{TP_{i}}{TP_{i}+FN_{i}}$ (28)
@@
[…]
@@ F.1 Additional Quantitative Analysis
F.3 Computational Efficiency

### v3

@@ Abstract
22footnotetext: Code and datasets are all available at the link: https://github.com/XPgogogo/scSiameseClu.
@@ 1 Introduction
Single-cell RNA sequencing (scRNA-seq) technology represents a significant advancement in bioinformatics, enabling the capture of comprehensive genetic information from individual cells Shapiro et al. (2013); Wang et al. (2025). Cell clustering, a key step in single-cell RNA sequencing analysis, groups cells by their gene expression patterns to uncover the complex characteristics of distinct cell populations and provide insights into their biological functions and interactions Kiselev et al. (2019); Yang et al. (2024). Moreover, clustering gene expression patterns further contributes to various downstream tasks, such as marker gene identification and cell type annotation Xu et al. (2025). Cluster analysis of scRNA-seq data has been a vibrant research area over the past decade.
In recent years, advanced computational methods have been increasingly explored to address the challenges of analyzing high-dimensional and sparse scRNA-seq data. Classical clustering methods, such as K-means, are straightforward and computationally efficient but struggle to capture the complex, nonlinear relationships in scRNA-seq data. Meanwhile, recent research has focused on applying deep learning frameworks for learning representations of scRNA-seq data and classifying samples into distinct clusters. A commonly used method is self-supervised learning methods, which uncover effective representations of scRNA-seq data by reconstructing the original input data Eraslan et al. (2019); Lopez et al. (2018); Tian et al. (2019). However, these methods primarily focus on extracting features from individual cells, overlooking the complex relationships between cells that are essential for understanding cellular diversity. Benefiting from the powerful utilization of graph information, graph neural networks (GNNs) have been applied to analyze scRNA-seq data by modeling cells as nodes and their interactions as edges, effectively capturing both gene expressions and cell graphs to comprehensively represent cellular heterogeneity  Wang et al. (2021b); Gan et al. (2022); Zhan et al. (2023).
@@ 2 Related Work
Deep Clustering for scRNA-seq. In scRNA-seq data analysis, early clustering algorithms like Phenograph Levine et al. (2015), MAGIC Van Dijk et al. (2018), and Seurat Wang et al. (2025) utilize k-nearest neighbor (KNN) graphs to model cell relationships, while SIMLR Wang et al. (2018) and MPSSC Park and Zhao (2018) employ multiple kernel functions to derive robust similarity measures from different data representations. These methods, however, often struggle with high-dimensional data and complex nonlinear features, making them susceptible to noise. Deep learning approaches have since become prominent. DCA Eraslan et al. (2019) employs a Zero-Inflated Negative Binomial (ZINB) autoencoder to model scRNA-seq data distributions, capturing nonlinear gene dependencies. SCVI Lopez et al. (2018) and SCVIS Ding et al. (2018), which rely on autoencoders for dimensionality reduction, often face over-regularization due to their Gaussian distribution assumption. Additionally, scDeepCluster Tian et al. (2019) combines ZINB-based autoencoders with deep embedding clustering to optimize feature learning and clustering. Most approaches ignore inter-gene and inter-cell correlations, focusing solely on individual cell expression profiles.
Graph Clustering for scRNA-seq. Recently, graph-based clustering methods have gained considerable attention for their effectiveness Tu et al. (2021); Liu et al. (2022); Ning et al. (2025). Methods like scGAE Luo et al. (2021) and graph-sc Ciortan and Defrance (2022) utilize graph autoencoders to embed scRNA-seq data while preserving topological structures. scGAC Cheng and Ma (2022) constructs cell graphs and employs a self-optimization approach for simultaneous representation learning and clustering. Notably, scGNN Wang et al. (2021b) uses GNNs and multi-modal autoencoders to aggregate cell-cell relationships and model gene expression patterns, while scDSC Gan et al. (2022) combines a ZINB model-based autoencoder with GNN modules using a mutually supervised strategy. These methods often struggle with representation collapse, severely limiting their ability to accurately model cell relationships.
@@ 3.2 Framework Overview
3.3 Dual Augmentation Module
@@
3.4 Siamese Fusion Module
@@
Latent Correlation Refinement. Similarly, latent correlation refinement (LCR) optimizes relationships between latent embeddings across different augmented views by aligning corresponding embeddings. First, we project gene expression embeddings $\mathbf{Z}^{G1}$ and $\mathbf{Z}^{G2}$ into $\widetilde{\mathbf{Z}}^{G1}$ and $\widetilde{\mathbf{Z}}^{G2}\in\mathbb{R}^{d\times K}$ with using a readout function $\mathcal{R}(\cdot):\mathbb{R}^{d\times N}\rightarrow\mathbb{R}^{d\times K}$ , formulated as $\widetilde{\mathbf{Z}}^{G1}=\mathcal{R}\left(\left(\mathbf{Z}^{G1}\right)^{\mathrm{T}}\right)$, $\widetilde{\mathbf{Z}}^{G2}=\mathcal{R}\left(\left(\mathbf{Z}^{G2}\right)^{\mathrm{T}}\right)$. Likewise, we can project cell graph embeddings $\mathbf{Z}^{C1}$ and $\mathbf{Z}^{C2}$ into $\widetilde{\mathbf{Z}}^{C1}=\mathcal{R}\left(\left(\mathbf{Z}^{C1}\right)^{\mathrm{T}}\right)$ and $\widetilde{\mathbf{Z}}^{C2}=\mathcal{R}\left(\left(\mathbf{Z}^{C2}\right)^{\mathrm{T}}\right)$.
@@
3.5 Optimal Transport Clustering
@@
In this context, we regard the target distribution $P$ as the transport plan matrix derived from optimal transport theory, while $\text{log}Q$ serves as the cost matrix. We impose the constraint $P^{T}\bm{1}_{N}=N\bm{\pi}$, where $\pi$ indicates the proportion of cells assigned to each cluster, estimated from intermediate clustering results. Thus, we can ensure that the resulting cluster distribution aligns with the defined mixing proportions. Given the substantial computational expense associated with direct optimization techniques, we utilize the Sinkhorn distance Sinkhorn (1967) to facilitate quicker optimization through an entropic constraint. The optimization problem, which includes a Lagrange multiplier for the entropy constraint is formulated as:
@@ 4 Experiments
In this section, we first validate our method through extensive experiments, demonstrating its superior performance in scRNA-seq clustering. We then show that the embeddings generated by our method effectively mitigate the embedding collapse issue in graph-based models. We also evaluate its utility in downstream tasks, like cell type annotation.
4.1 Experimental Setup
@@
Evaluation Metrics. To illustrate the effectiveness of our approach, we evaluate clustering performance through three widely recognized metrics: Accuracy (ACC), Normalized Mutual Information (NMI) Strehl and Ghosh (2002), and Adjusted Rand Index (ARI) Vinh et al. (2009).
@@ 5 Conclusion
Acknowledgements
This work is partially supported by the National Natural Science Foundation of China (Grant No. 92470204 and 62406306), the National Key Research and Development Program of China (Grant No. 2024YFF0729201), and the Science and Technology Development Fund (FDCT), Macau SAR (file no. 0123/2023/RIA2, 001/2024/SKL).
@@ C.3 Gene expression and Cell Graph Autoencoders
$\begin{aligned} \mathcal{L}_{REC}&=\mathcal{L}_{GEA}+\mathcal{L}_{CGA}\\ &=\frac{1}{2N}\|\widetilde{\mathbf{X}}-\widehat{\mathbf{X}}\|_{F}^{2}+\left(\frac{1}{2N}\|\widetilde{\mathbf{A}}\widetilde{\mathbf{X}}-\mathbf{Z}\|_{F}^{2}+\eta\left(\frac{1}{2N}\|\widetilde{\mathbf{A}}-\widehat{\mathbf{A}}\|_{F}^{2}\right)\right),\end{aligned}$
@@
$\begin{aligned} \mathrm{NB}(\mathbf{X}_{ij};\mu_{ij},\theta_{ij})=\frac{\Gamma(\mathbf{X}_{ij}+\theta_{ij})}{\Gamma(\theta_{ij})}\left(\frac{\theta_{ij}}{\theta_{ij}+\mu_{ij}}\right)^{\theta_{ij}}\left(\frac{\mu_{ij}}{\theta_{ij}+\mu_{ij}}\right)^{\mathbf{X}_{ij}},\end{aligned}$
@@ D.1 Evaluation Metrics
$\text{Macro Precision}=\frac{1}{N}\sum^{N}_{i=1}\frac{TP_{i}}{TP_{i}+FP_{i}}$ (27)
@@
$\text{Macro Recall}=\frac{1}{N}\sum^{N}_{i=1}\frac{TP_{i}}{TP_{i}+FN_{i}}$ (28)
@@
$\text{Macro F1}=\frac{1}{N}\sum^{N}_{i=1}\frac{2*Precision_{i}*Recall_{i}}{Precision_{i}+Recall_{i}}$ (29)
@@ F.1 Additional Quantitative Analysis
F.3 Ablation Study
(a) Major components evaluaion
(b) Component Effects in SFM
Figure 10: Ablation study on  Shekhar mouse retina cells dataset.
F.4 Computational Efficiency